# A Stable Whitening Optimizer for Efficient Neural Network Training

### v2

@@ 6 Discussion and Conclusion
Empirically, we show that SPlus can achieve the same validation performance as Adam with $\sim 44\%$ of the gradient steps, and in $\sim 62\%$ of the wall-clock time. Over a range of training objectives and checkpoints, SPlus achieves the state-of-the-art performance in comparison to previous optimizers.
@@
Limitations. Computationally, SPlus requires storing $3nm+2(n^{2}+m^{2})$ parameters per dense layer–three instances for the live parameters, slow parameters, and momentum, along with two Kronecker factors for the gradient covariances plus the cached eigenvectors. For a square matrix, this is roughly $60\%$ more memory than Adam. Furthermore, SPlus requires additional wall-clock time – there is an additional matrix multiplication in each update to rotate via the eigenbasis (practically, this is negligible), and a more significant wall-clock cost for matrix eigendecomposition. The training settings in this work only consider Transformer architectures. While this is an intentional choice, as neural networks today have largely converged on the Transformer backbone, it remains unanswered how performance would vary on non-Transformer architectures.
@@
7 How should I use SPlus as a practitioner?
We have done our best to make swapping in SPlus as easy as possible. We provide single-file implementations of SPlus in both JAX and Pytorch at the repo github.com/kvfrans/splus. Check the README for additional details. In general, to replace a tuned Adam with SPlus, we recommend the following formula for learning rate conversion:
[…]
or, as a reasonable starting point,
[…]
Other common hyperparameters ($\beta_{1}$, $\beta_{2}$, weight decay) can generally be kept unchanged. We found that SPlus-specific hyperparameters (EMA rate, constant factor for nonstandard parameters, inversion frequency) are relatively unsensitive, and do not need to be tuned from their default values.
Importantly, SPlus requires evaluating at a different set of parameters than used for gradient calculation, which unfortunately does not fit nicely with mainstream training abstractions. We provide helper functions in the JAX/Pytorch implementations to assist with this.
@@ A.1 Pseudocode of SPlus
{minted}
python class SPlusState(NamedTuple): ema: chex.Array momentum: chex.Array sides: chex.Array q_sides: chex.Array step: int ema_rate: float
def splus_get_eval_params(state): ema_hat = jax.tree_map(lambda e: e / (1 - state.ema_rate ** state.step), state.ema) return ema_hat
def splus( learning_rate: base.ScalarOrSchedule, b1: float = 0.9, b2: float = 0.999, ema_rate: float = 0.999, eps: float = 1e-30, inverse_every: int = 100, nonstandard_constant: float = 0.001, weight_decay: float = 1e-2, mask: Optional[Union[Any, Callable[[base.Params], Any]]] = None, max_dim: int = 10000, ):
def init_fn(params): momentum = otu.tree_zeros_like(params) ema = otu.tree_zeros_like(params) def sides_decomp(p): if len(p.shape) == 2: return [jnp.zeros((d, d)) if d < max_dim else None for d in p.shape] return None sides = jax.tree_map(sides_decomp, params) def qs_decomp(p): if len(p.shape) == 2: return [jnp.eye(d) if d < max_dim else None for d in p.shape] q_sides = jax.tree_map(qs_decomp, params) step = 0 return SPlusState(ema, momentum, sides, q_sides, step, ema_rate)
def update_sides(g, s): if len(g.shape) == 2: return [ b2 * s[0] + (1 - b2) * g @ g.T if s[0] is not None else None, b2 * s[1] + (1 - b2) * g.T @ g if s[1] is not None else None, ] else: return None
def rot(p, q): if len(p.shape) == 2: p = q[0].T @ p if q[0] is not None else p p = p @ q[1] if q[1] is not None else p return p
def unrot(p, q): if len(p.shape) == 2: p = q[0] @ p if q[0] is not None else p p = p @ q[1].T if q[1] is not None else p return p
@jax.jit def get_eigvecs(s): if s is None: return None _, q = jnp.linalg.eigh(s + eps * jnp.eye(s.shape[0])) return q
def update_inverse(sides): q_sides = jax.tree_map(get_eigvecs, sides) return q_sides
def update_fn(grads, state, params): step = state.step + 1
# Rotate to eigenbasis, take sign, unrotate. momentum = jax.tree_map(lambda m, g: b1 * m + (1 - b1) * g, state.momentum, grads) momentum_rot = jax.tree_map(rot, momentum, state.q_sides) updates_rot = jax.tree_map(lambda m: jnp.sign(m), momentum_rot) updates = jax.tree_map(unrot, updates_rot, state.q_sides) sides = jax.tree_map(update_sides, grads, state.sides) ema = jax.tree_map(lambda e, g: ema_rate * e + (1 - ema_rate) * g, state.ema, params)
# Every ‘inverse_every‘ steps, we update the inverse eigendecomposition. do_inverse = (step q_sides = jax.lax.cond(do_inverse, update_inverse, lambda _: state.q_sides, sides)
return updates, SPlusState(ema, momentum, sides, q_sides, step, state.ema_rate)
def shape_scaling(updates, state, params): def shape_scale(path, u): path_str = ’/’.join([p.key for p in path]) if len(u.shape) == 2 and u.shape[0] < max_dim and u.shape[1] < max_dim: scale = (1 / (u.shape[0] + u.shape[1])/2) else: scale = nonstandard_constant return u * scale return jax.tree_util.tree_map_with_path(shape_scale, updates), None
splus_main = base.GradientTransformation(init_fn, update_fn) splus_scaling = base.GradientTransformation(lambda _: None, shape_scaling) return combine.chain( splus_main, transform.add_decayed_weights(weight_decay, mask), transform.scale_by_learning_rate(learning_rate), splus_scaling )
@@ A.2 Full results of optimizer comparisons
We present in Table 1 and Table 2 an extended table of results, comparing optimizer performance under each specific objective and starting checkpoint. As described in Section 5, all experiments are conducted using a 160M parameter Transformer model. In each plot, we compare the amount of gradient steps and/or wallclock time required to match the performance of Adam. All optimizers are trained for 10k gradient steps, and the learning rate is tuned independently. "Div." indicates that training diverges under any non-trivial learning rate. "$>1.0$" indicates that at the 10k step mark, the method is unable to outperform Adam.
@@ A.3 Spectral scaling vs symmetric scaling
In Section 4.2, we mention a difference between the SPlus symmetric scaling factor $2/(m+n)$ versus the "spectral" scaling [54] which argues for $1/m$. The spectral scaling is the correct factor such that regardless of any dense layer input/output ratio, the scale of activation updates remains constant. However, we find experimentally that this property is harmful for transformer training. One hypothesis is that for non-square dense layers, gradients are effectively low-rank and/or certain eigenbases do not align with incoming activation vectors.
As shown in Figure 7, the symmetric scaling that we opt for maintains the original performance, outperforming spectral scaling while allowing for learning rate transfer. The rationale behind symmetric scaling is that for the core Transformer backbones – specifically the MLP block which often consists of two layers of $(m,4m)$ and $(4m,m)$ size – learning follows the same trajectory as if no scaling factor was used at all, i.e. a single global learning rate is applied to all parameters.
@@
$u=\operatorname*{arg\,min}_{\Delta\theta}\;\underbrace{\;g^{T}\Delta\theta\;}_{\text{Improvement}}+\underbrace{\Delta\theta^{T}M\Delta\theta}_{\text{Distance Penalty}}\quad=\quad M^{-1}g.$ (19)
@@
Note the expectation in Equation 23, which notes that $y$ must be sampled from the current distribution. This means that the Fisher cannot be calculated by taking the loss over samples from the dataset, and must instead use sampled outputs.
@@
where notably the second-order terms in the form of Equation 21 evaluate to zero.

### v3

@@ 6 Discussion and Conclusion
Empirically, we show that SPlus can achieve the same validation performance as Adam with $\sim 44$-$58\%$ of the gradient steps, and in $\sim 62$-$83\%$ of the wall-clock time. Over a range of training objectives and checkpoints, SPlus achieves superior performance in comparison to previous optimizers.
@@
Limitations. In comparison to Adam, SPlus requires storing $3nm+2(n^{2}+m^{2})$ parameters per dense layer–three instances for the live parameters, slow parameters, and momentum, along with two Kronecker factors for the gradient covariances plus the cached eigenvectors. For a square matrix, this is roughly $60\%$ more memory than Adam. Furthermore, SPlus requires nontrivial wall-clock time for matrix eigendecomposition. We note that SPlus (without iterate-averaging) uses the same amount of compute as Shampoo, and less memory than SOAP. The training settings in this work only consider Transformer architectures. While this is an intentional choice, as neural networks today have largely converged on the Transformer backbone, it remains unanswered how performance would vary on non-Transformer architectures. We utilize a constant learning rate as is common in diffusion model training [37], but is less standard in language model training.
@@ A.1 Pseudocode of SPlus
class SPlusState(NamedTuple):
ema: chex.Array
momentum: chex.Array
sides: chex.Array
q_sides: chex.Array
step: int
ema_rate: float
[…]
ema_hat = jax.tree_map(lambda e: e / (1 - state.ema_rate ** state.step), state.ema)
return ema_hat
[…]
learning_rate: base.ScalarOrSchedule,
b1: float = 0.9,
b2: float = 0.999,
ema_rate: float = 0.999,
eps: float = 1e-30,
inverse_every: int = 100,
nonstandard_constant: float = 0.001,
weight_decay: float = 1e-2,
mask: Optional[Union[Any, Callable[[base.Params], Any]]] = None,
max_dim: int = 10000,
[…]
momentum = otu.tree_zeros_like(params)
ema = otu.tree_zeros_like(params)
def sides_decomp(p):
if len(p.shape) == 2:
return [jnp.zeros((d, d)) if d < max_dim else None for d in p.shape]
return None
sides = jax.tree_map(sides_decomp, params)
def qs_decomp(p):
if len(p.shape) == 2:
return [jnp.eye(d) if d < max_dim else None for d in p.shape]
q_sides = jax.tree_map(qs_decomp, params)
step = 0
return SPlusState(ema, momentum, sides, q_sides, step, ema_rate)
[…]
if len(g.shape) == 2:
return [
b2 * s[0] + (1 - b2) * g @ g.T if s[0] is not None else None,
b2 * s[1] + (1 - b2) * g.T @ g if s[1] is not None else None,
else:
return None
[…]
if len(p.shape) == 2:
p = q[0].T @ p if q[0] is not None else p
p = p @ q[1] if q[1] is not None else p
return p
[…]
if len(p.shape) == 2:
p = q[0] @ p if q[0] is not None else p
p = p @ q[1].T if q[1] is not None else p
return p
[…]
def get_eigvecs(s):
if s is None:
return None
_, q = jnp.linalg.eigh(s + eps * jnp.eye(s.shape[0]))
return q
[…]
q_sides = jax.tree_map(get_eigvecs, sides)
return q_sides
[…]
step = state.step + 1
[…]
momentum = jax.tree_map(lambda m, g: b1 * m + (1 - b1) * g, state.momentum, grads)
momentum_rot = jax.tree_map(rot, momentum, state.q_sides)
updates_rot = jax.tree_map(lambda m: jnp.sign(m), momentum_rot)
updates = jax.tree_map(unrot, updates_rot, state.q_sides)
sides = jax.tree_map(update_sides, grads, state.sides)
ema = jax.tree_map(lambda e, g: ema_rate * e + (1 - ema_rate) * g, state.ema, params)
[…]
do_inverse = (step q_sides = jax.lax.cond(do_inverse, update_inverse, lambda _ : state.q_sides, sides)
[…]
def shape_scale(path, u):
path_str = ’/’.join([p.key for p in path])
if len(u.shape) == 2 and u.shape[0] < max_dim and u.shape[1] < max_dim:
scale = (1 / (u.shape[0] + u.shape[1])/2)
else:
scale = nonstandard_constant
return u * scale
return jax.tree_util.tree_map_with_path(shape_scale, updates), None
[…]
splus_scaling = base.GradientTransformation(lambda _ : None, shape_scaling)
return combine.chain(
splus_main,
transform.add_decayed_weights(weight_decay, mask),
transform.scale_by_learning_rate(learning_rate),
splus_scaling
@@ A.2 Full results of optimizer comparisons
We present in Table˜3 and Table˜4 an extended table of results, comparing optimizer performance under each specific objective and starting checkpoint. As described in Section˜5, all experiments are conducted using a 160M parameter Transformer model. In each plot, we compare the amount of gradient steps and/or wallclock time required to match the performance of Adam. All optimizers are trained for 10k gradient steps, and the learning rate is tuned independently. "Div." indicates that training diverges under any non-trivial learning rate. ">1.0>1.0" indicates that at the 10k step mark, the method is unable to outperform Adam.
@@ A.3 Spectral scaling vs symmetric scaling
In Section˜4.2, we mention a difference between the SPlus symmetric scaling factor $2/(m+n)$ versus the "spectral" scaling [55] which argues for $1/m$. The spectral scaling is the correct factor such that regardless of any dense layer input/output ratio, the scale of activation updates remains constant. However, we find experimentally that this property is harmful for transformer training. One hypothesis is that for non-square dense layers, gradients are effectively low-rank and/or certain eigenbases do not align with incoming activation vectors.
As shown in Figure˜7, the symmetric scaling that we opt for maintains the original performance, outperforming spectral scaling while allowing for learning rate transfer. The rationale behind symmetric scaling is that for the core Transformer backbones – specifically the MLP block which often consists of two layers of $(m,4m)$ and $(4m,m)$ size – learning follows the same trajectory as if no scaling factor was used at all, i.e. a single global learning rate is applied to all parameters.
@@
$u=\operatorname*{arg\,min}_{\Delta\theta}\;\underbrace{\;g^{T}\Delta\theta\;}_{\text{Improvement}}+\underbrace{(1/2)\Delta\theta^{T}M\Delta\theta}_{\text{Distance Penalty}}\quad=\quad M^{-1}g.$ (17)
@@
Note the expectation in Equation˜21, which notes that $y$ must be sampled from the current distribution. This means that the Fisher cannot be calculated by taking the loss over samples from the dataset, and must instead use sampled outputs.
@@
where notably the second-order terms in the form of Equation˜19 evaluate to zero.
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We describe the empirical setting, then show the according results in the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We include a limitations section in the final page.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not provide theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have put forth a best-faith effort to provide reproducible training details and methodology in section 5. Nuanced specifics can be found within the provided code.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We only utilize open-source datasets in this work, and provide access to our code.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have put forth a best-faith effort to provide reproducible training details and methodology in section 5. Nuanced specifics can be found within the provided code.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Due to computational constraints, we are only able to run one trial each for each optimizer, under a specific objective, checkpoint, and learning rate. That said, we control for random seed and data order, and heuristically the training runs do not large independent variance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We describe the machines used to run the experiments, and give an estimate of the time taken.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We follow the ethics guidelines.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: We do not find the need to discuss specific societal impacts of this work, as it purely relates to a computational problem.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We propose a computational method, and do not release any new data.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the two open-source datasets that we utilize.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not involve crowdsourcing or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.